Cricket World Cup (CWC) is considering a project proposal which requires an initial investment of $72,625 and it is expected to have net cash flows of $15,000 per year for 8 years. The firm cash flows are discounted at a rate of 12 percent.

a. What is the project’s Net Present Value (NPV)? (Rounded to 2 decimal places)

b. What is the project’s discounted payback period? (Rounded to 2 decimal places


1
Expert's answer
2021-04-19T18:50:23-0400

a. The project’s Net Present Value (NPV) is:

"NPV = - 72,625 + 15,000\/1.12 + 15,000\/1.12^2 + 15,000\/1.12^3 + 15,000\/1.12^4 + 15,000\/1.12^5 + 15,000\/1.12^6 + 15,000\/1.12^7 + 15,000\/1.12^8 = 1889.60."

b. What is the project’s discounted payback period is:

"PBP = 7 + (1 - 1,889.6\/(15,000\/1.12^8)) = 7.69" years.